# On geometric spectral functionals

### v1

@@
On geometric spectral functionals
@@
Arkadiusz Bochniak a, Ludwik Dąbrowski b, Andrzej Sitarz c and Paweł Zalecki c
@@
A. Bochniak, L. Dabrowski, A. Sitarz and P. Zalecki
@@
a) Max-Planck-Institut für Quantenoptik, Hans-Kopfermann-Str. 1, Garching, 85748, Germany;
@@
1 Introduction
Methods from Riemannian and Lorentzian geometry provide the mathematical foundation for modern theories of physical interactions. These differential geometric frameworks are central not only to Einstein’s General Theory of Relativity — which has been extensively validated through observations — but also to a broad array of theoretical extensions and modifications that seek to address unresolved questions in cosmology and explore phenomena beyond the classical scope of General Relativity. Differential operators defined on a smooth manifold play a fundamental role in the mathematical formulation of physical theories and are key elements in both quantum mechanics and quantum field theory. Their spectral properties determine the energy levels of systems and influence many physical phenomena.
From a mathematical perspective, the study of spectra of such operators — such as the Laplace-Beltrami operator $\Delta$ or the Dirac operator $D$ — has led to the emergence and rapid development of a vibrant branch of mathematics known as spectral geometry. This field investigates how the eigenvalues and eigenfunctions of differential operators reflect the geometric and topological features of the underlying space. One of the most iconic and influential questions in this context was popularized by Mark Kac: “Can one hear the shape of a drum?[14]. Though the answer is nuanced and context-dependent, it inspired a deeper exploration of the relationship between geometry and analysis.
This line of inquiry has since evolved into a powerful paradigm, suggesting that key properties of physical systems — and even the structure of spacetime itself — can be encoded in the spectrum of certain operators. In particular, within the framework of Alain Connes’ noncommutative geometry [5, 4] and spectral geometry, this idea has been extended to propose that the dynamics and topology of spacetime may be recovered from the spectral data of appropriate differential operators. This approach offers a compelling bridge between abstract mathematical structures and physical reality.
In classical spectral geometry, the primary objects of study are spectral quantities defined for (pseudo-)differential operators acting on sections of vector bundles over smooth manifolds. They typically depend on the asymptotics of the spectrum of these operators and often take the form of exotic traces, heat kernel coefficients, zeta functions, or determinant-like quantities.
When these operators act on sections of vector bundles — such as the tangent bundle or the spinor bundle — they reflect additional geometric and topological structures. For example, the heat trace asymptotics link the short-time behavior of the heat kernel to local geometric invariants like scalar curvature, while the spectral zeta functions encode global properties [10, 11].
An alternative and particularly elegant approach to defining spectral functionals makes use of the Wodzicki residue $\mathscr{W}\!\mathrm{res}$ — a powerful tool in the analysis of pseudo-differential operators. The Wodzicki residue is remarkable in that it provides the only trace, up to multiplication by a constant, on the algebra of classical pseudo-differential operators acting on sections of a complex vector bundle over a compact, oriented manifold of dimension $n\geq 2$ [18, 12]. It extends the classical notion of a trace to a broader class of operators by assigning a well-defined density that captures geometric information, in particular linking the local structure of the operator to global geometric invariants. The local densities appearing in spectral functionals of geometric significance correspond to geometric invariants of the underlying manifold or vector bundle, thereby justifying the nomenclature adopted for such functionals. These functionals can then be generalized to more abstract spectral triples within the framework of noncommutative geometry.
Let us briefly review the spectral functionals of particular physical interest. First, recall that the metric functional $\mathscr{g}_{D}$ and the Einstein functional $\mathscr{G}_{D}$ were constructed for spectral triples [7, Definition 5.4] and, respectively, assign to a pair of one-forms $(u,w)$ a number:
$\mathscr{g}_{D}(u,w)=\mathscr{W}\!\mathrm{res}(\hat{u}\hat{w}|D|^{-n}),\qquad u,w\in\Omega_{D}^{1},$ (1.1)
and
$\mathscr{G}_{D}(u,w)=\mathscr{W}\!\mathrm{res}\bigl{(}\hat{u}{D,\hat{w}}D|D|^{-n}\bigr{)},\qquad u,w\in\Omega_{D}^{1},$ (1.2)
The torsion functional, as defined in [9, Definition 2.1], assigns to a triple of one-forms $(u,v,w)$:
$\mathscr{T}_{D}(u,v,w)=\mathscr{W}\!\mathrm{res}(\hat{u}\hat{v}\hat{w}DD^{-2m}),\qquad u,v,w\in\Omega^{1}_{D}.$ (1.3)
Finally, the scalar curvature functional is defined by $\mathscr{R}_{D}(f)=\mathscr{W}\!\mathrm{res}(fD^{-2m+2})$ for $f\in\mathcal{A}$ [15, 1]. This functional, when regarded as a functional depending on the metric that determines the Dirac operator, is commonly known as the Einstein–Hilbert functional, since it directly corresponds to the Einstein–Hilbert action — the cornerstone of General Relativity.
2 Spectral functionals: preliminaries
In this section, we first recall the results of computations of Wodzicki residues for a broad class of Laplace- and Dirac-type operators acting on sections of a vector bundle $V$ over a manifold $M$. Here, $M$ always denotes a compact, closed Riemannian manifold of dimension $n=2m>1$ with a fixed metric [^1] We remark that most of the results presented here remain valid in the odd-dimensional case, as observed in [3] for the spin-Dirac operator. For the sake of concreteness and generalizations involving gradings, we focus here on even-dimensional manifolds..
We follow the conventions and notation used previously in [7, 3]; for example, we work in normal coordinates at a fixed point and expand the homogeneous symbols of the operators under consideration in these coordinates. We recall the notion of a Laplace-type operator discussed in [3]. An operator $L$ acting on $\Gamma(V)$, the module of smooth sections of $V$, is said to be of Laplace type if its symbol at any point $x\in M$ is given by $\sigma(L)=\mathfrak{a}_{2}+\mathfrak{a}_{1}+\mathfrak{a}_{0}$, where the homogeneous symbols $\mathfrak{a}_{\bullet}$, computed in normal coordinates $x$, are:
$\begin{split}\mathfrak{a}_{2}&=\bigl{(}\delta_{ab}+\frac{1}{3}R_{acbd}x^{c}x^{d}\bigr{)}\xi_{a}\xi_{b}+o({\bf x}^{2}),\\ \mathfrak{a}_{1}&=i(P_{ab}x^{b}+S_{a})\xi_{a}+o({\bf x}),\\ \mathfrak{a}_{0}&=Q+o({\bf 1})\end{split}$ (2.1)
Here, $R_{acbd}$ is the Riemann curvature tensor, and $P_{ab}$, $S_{a}$, and $Q$ are $\mathrm{End}(V)$-valued tensors evaluated at $x=0$. We recall that the Wodzicki residue [18] is (up to normalization) the unique trace on the algebra $\Psi\mathrm{DO}(V)$ of classical pseudodifferential operators:
$\mathscr{W}\!\mathrm{res}(P)=\int_{M}\mathrm{dvol}_{g}\int_{|\xi|=1}\!\!\!\!d^{4}x,\mathrm{Tr}\bigl{(}\sigma_{-n}(P)(x,\xi)\bigr{)},$ (2.2)
where $\sigma_{-n}(P)$ denotes the symbol of $P\in\Psi\mathrm{DO}(V)$ of order $-n$ [10].
To compute the relevant functionals, we need the following general result:
Proposition 2.1.
[3, Proposition I.2] For a second-order differential operator $O$, with a symbol expressed in normal coordinates around a point on $M$,
$\sigma(O)=F^{ab}\xi_{a}\xi_{b}+iG^{a}\xi_{a}+H+o({\bf 1}),$
where $F^{ab}=F^{ba}$, $G^{a}$, and $H$ are endomorphisms of the fiber $V$ at the point $x=0$, we have:
$\mathscr{W}\!\mathrm{res}(OL^{-m})=\frac{\nu_{n-1}}{24}$ $\int_{M}\mathrm{dvol}_{g}\mathrm{Tr}\,\bigl{[}24H+12G^{a}S_{a}+F^{aa}(-12Q+6P_{bb}-2R-3S_{b}S_{b})$ (2.3)
$\hskip 56.9055pt+2F^{ab}(-6P_{ab}+2\mathrm{Ric}_{ab}-3S_{a}S_{b})\bigr{]},$
where $R$ is the scalar curvature and $\mathrm{Ric}$ is the Ricci tensor.
We now present two corollaries that will be particularly useful.
Corollary 2.2.
[3, Corollary I.3] For a $C^{\infty}(M)$–endomorphism $E:S\to S$ we have:
$\mathscr{W}\!\mathrm{res}(EL^{-m+1})=\frac{n-2}{24}\nu_{n-1}\int_{M}\mathrm{dvol}_{g}\mathrm{Tr}\left[E\left(-12Q+6P_{aa}-2R-3S_{a}S_{a}\right)\right].$ (2.4)
Corollary 2.3.
[3, Corollary I.4] For an operator $O$ with a symbol as in Proposition 2.1 we have,
$\mathscr{W}\!\mathrm{res}\Biggl{(}\Bigl{(}O-\frac{1}{n-2}F^{aa}L\Bigr{)}L^{-m}\Biggr{)}$ $=\frac{\nu_{n-1}}{24}\int_{M}\mathrm{dvol}_{g}\mathrm{Tr}[24H+12G^{a}S_{a}$ (2.5)
$\phantom{xxxx}+2F^{ab}(-6P_{ab}+2\mathrm{Ric}_{ab}-3S_{a}S_{b})].$
We impose minimal assumptions on $D_{0}$ ensuring that $D_{0}^{2}$ is of Laplace type:
Definition 2.4.
We say that $D=D_{0}+B$ is a Dirac-type operator on sections of the vector bundle $V$ if the symbol of $D_{0}$ in normal coordinates around a fixed point on the manifold is given by
$\sigma(D_{0})=\mathfrak{d}_{1}+\mathfrak{d}_{0},$
where
$\mathfrak{d}_{1}=(-\gamma^{a}+f^{a}_{bc}x^{b}x^{c})\xi_{a}+o({\bf x^{2}}),\qquad\mathfrak{d}_{0}=g_{a}x^{a}+o({\bf x}),$
for some endomorphisms $f^{a}_{bc}$ (symmetric in $b,c$) and $g_{a}$ of $V$ evaluated at $x=0$. Additionally, the endomorphism $B$ has an expansion:
$B=B_{0}+B_{a}x^{a}+o({\bf x}).$
The structure of $\mathfrak{d}_{1}$ and $\mathfrak{d}_{0}$, together with the requirement that $\mathfrak{d}_{1}^{2}=\mathfrak{a}_{2}$ as in (2.1), ensures that both $D_{0}^{2}$ and $D^{2}$ are Laplace-type. Notably, the absence of linear terms in ${\bf x}$ in the principal symbol $\mathfrak{d}_{1}$, and the vanishing of the zero-order symbol $\mathfrak{d}_{0}$ at $x=0$, are essential conditions [^2] One could define Dirac-type operators without requiring the absence of linear terms in ${\bf x}$ in the principal symbol. However, for our purposes, this condition is essential.. The explicit form of the symbols of $D_{0}$ (e.g., coefficients $f^{a}_{bc}$ and $g_{a}$) will not matter, as we focus only on the perturbation by $B$. The $\gamma^{a}$ are Clifford algebra generators on each fiber $V$ satisfying ${\gamma^{a},\gamma^{b}}=2\delta^{ab}\mathbf{1}$. Finally, only the first two terms in the expansion of $B$ are relevant, since $B$ does not enter the principal symbol of $D$.
Proposition 2.5.
Let $D$ be as in Definition 2.4. Then, for $O=ED$, where $E$ is an arbitrary bundle endomorphism, we have:
$\mathscr{W}\!\mathrm{res}(EDD^{-2m})=\frac{\nu_{n-1}}{2}\int_{M}\mathrm{dvol}_{g}\mathrm{Tr}\left[E(2B_{0}-\gamma^{a}\{\gamma^{a},B_{0}\})\right].$ (2.6)
Proof 2.6.
Observe that $O=ED$ has the structure from Proposition 2.1, with $F^{ab}=0$, $G^{a}=iE\gamma^{a}$, and $H=EB_{0}$. The result follows directly from Proposition 2.1, with $L=D^{2}$, and the following observations: For $k>0$, the leading symbols of $D^{-2k}$ read
$\sigma(D^{-2k})=\mathfrak{c}_{2k}+\mathfrak{c}_{2k+1}+\mathfrak{c}_{2k+2},$
where
$\mathfrak{c}_{2k}=\mathfrak{c}^{0}_{2k}+o({\bf x^{2}}),$ (2.7)
$\mathfrak{c}_{2k+1}=\mathfrak{c}^{0}_{2k+1}+k\xi_{a}||\xi||^{-2k-2}\{\gamma^{a},B\}+o({\bf x}),$
$\mathfrak{c}_{2k+2}=\mathfrak{c}^{0}_{2k+2}-k||\xi||^{-2k-2}(i\gamma^{a}B_{a}+B^{2}_{0})$
$\qquad\qquad+k(k+1)||\xi||^{-2k-4}\left(i\{\gamma^{a},B_{b}\}+\frac{1}{2}\{\gamma^{a},B_{0}\}\{\gamma^{b},B_{0}\}\right)\xi_{a}\xi_{b}+o({\bf 1}).$
Here, $\mathfrak{c}_{p}^{0}$ denotes the respective symbols of $D_{0}^{-2k}$:
$\sigma(D_{0}^{-2k})=\mathfrak{c}_{2k}^{0}+\mathfrak{c}_{2k+1}^{0}+\mathfrak{c}_{2k+2}^{0}.$
Since $L=D^{2}=D_{0}^{2}+\{D_{0},B\}+B^{2}$, and using the relation $\{i\gamma^{a}\partial_{a},B\}=i\{\gamma^{a},B\}\partial_{a}+i\gamma^{a}\partial_{a}B$, we get:
$P_{ab}=P_{ab}^{0}+i{\gamma^{a},B_{b}},\quad S_{a}=i{\gamma^{a},B_{0}},$ (2.8)
$S_{a}^{0}=0,\quad Q=Q^{0}+i\gamma^{a}B_{a}+B_{0}^{2},$
from which the claim follows.

### v2

@@
On Geometric Spectral FunctionalsThis paper is a contribution to the Special Issue on Asymptotics, Randomness and Noncommutativity. The full collection is available at https://sigma-journal.com/noncommutativity.html
@@
Arkadiusz BOCHNIAK ab, Ludwik DA̧BROWSKI c, Andrzej SITARZ d and Paweł ZALECKI d
@@
A. Bochniak, L. Da̧browski, A. Sitarz and P. Zalecki
@@
a) Max-Planck-Institut für Quantenoptik, Hans-Kopfermann-Str. 1, Garching, 85748, Germany
[…]
b) Munich Center for Quantum Science and Technology,
@@
d) Institute of Theoretical Physics, Jagiellonian University,
[…]
Received May 23, 2025, in final form March 24, 2026; Published online April 14, 2026
[…]
We investigate spectral functionals associated with Dirac and Laplace-type differential operators on manifolds, defined via the Wodzicki residue, extending classical results for Dirac operators derived from the Levi-Civita connection to geometries with torsion. The local densities of these functionals recover fundamental geometric tensors, including the volume form, Riemannian metric, scalar curvature, Einstein tensor, and torsion tensor. Additionally, we introduce chiral spectral functionals using a grading operator, which yields novel spectral invariants. These constructions offer a richer spectral-geometric characterization of manifolds.
[…]
spectral geometry; Wodzicki residue; noncommutative geometry; torsion
[…]
58B34; 58J50; 46L87; 83C65
1 Introduction
Methods from Riemannian and Lorentzian geometry provide the mathematical foundation for modern theories of physical interactions. These differential geometric frameworks are central not only to Einstein’s general theory of relativity – which has been extensively validated through observations – but also to a broad array of theoretical extensions and modifications that seek to address unresolved questions in cosmology and explore phenomena beyond the classical scope of general relativity. Differential operators defined on a smooth manifold play a fundamental role in the mathematical formulation of physical theories and are key elements in both quantum mechanics and quantum field theory. Their spectral properties determine the energy levels of systems and influence many physical phenomena.
From a mathematical perspective, the study of spectra of such operators – such as the Laplace–Beltrami operator $\Delta$ or the Dirac operator $D$ – has led to the emergence and rapid development of a vibrant branch of mathematics known as spectral geometry. This field investigates how the eigenvalues and eigenfunctions of differential operators reflect the geometric and topological features of the underlying space. One of the most iconic and influential questions in this context was popularized by Mark Kac: “Can one hear the shape of a drum?[15]. Though the answer is nuanced and context-dependent, it inspired a deeper exploration of the relationship between geometry and analysis.
This line of inquiry has since evolved into a powerful paradigm, suggesting that key properties of physical systems – and even the structure of spacetime itself – can be encoded in the spectrum of certain operators. In particular, within the framework of Alain Connes’ noncommutative geometry [4, 5] and spectral geometry, this idea has been extended to propose that the dynamics and topology of spacetime may be recovered from the spectral data of appropriate differential operators. This approach offers a compelling bridge between abstract mathematical structures and physical reality.
In classical spectral geometry, the primary objects of study are spectral quantities defined for (pseudo-)differential operators acting on sections of vector bundles over smooth real manifolds. They typically depend on the asymptotics of the spectrum of these operators and often take the form of exotic traces, heat kernel coefficients, zeta functions, or determinant-like quantities.
When these operators act on sections of vector bundles – such as the tangent bundle or the spinor bundle – they reflect additional geometric and topological structures. For example, the heat trace asymptotics link the short-time behavior of the heat kernel to local geometric invariants like scalar curvature, while the spectral zeta functions encode global properties [10, 11].
An alternative and particularly elegant approach to defining spectral functionals makes use of the Wodzicki residue $\mathscr{W}\mathrm{res}$ – a powerful tool in the analysis of pseudo-differential operators. The Wodzicki residue is remarkable in that it provides, up to multiplication by a constant, the unique trace on the space of classical pseudodifferential operators acting on sections of a complex vector bundle over a compact, oriented manifold of dimension $n\geq 2$, when considered modulo smoothing operators and within fixed integer orders [12, 19]. Rather than extending the usual operator trace, it defines a trace functional on a broader class of operators as the integral of a well-defined local density, thereby encoding geometric information and relating the local structure of an operator to global geometric invariants. The local densities appearing in spectral functionals of geometric significance correspond to geometric invariants of the underlying manifold or vector bundle, thereby justifying the nomenclature adopted for such functionals. These functionals can then be generalized to more abstract spectral triples within the framework of noncommutative geometry.
Let us briefly review the spectral functionals of particular physical interest. First, recall that the metric functional $\mathscr{g}_{D}$ and the Einstein functional $\mathscr{g}_{D}$ were constructed for spectral triples [7, Definition 5.4] and, respectively, assign to a pair of one-forms $(u,w)$ a number
$\mathscr{g}_{D}(u,w)=\mathscr{W}\mathrm{res}(\hat{u}\hat{w}|D|^{-n}),\qquad u,w\in\Omega_{D}^{1},$
and
[…]
The torsion functional, as defined in [8, Definition 2.1], assigns to a triple of one-forms $(u,v,w)$
$\mathscr{T}_{D}(u,v,w)=\mathscr{W}\mathrm{res}\bigl(\hat{u}\hat{v}\hat{w}DD^{-2m}\bigr),\qquad u,v,w\in\Omega^{1}_{D}.$
Finally, the scalar curvature functional is defined by $\mathscr{R}_{D}(f)=\mathscr{W}\mathrm{res}\bigl(fD^{-2m+2}\bigr)$ for $f\in\mathcal{A}$ [1, 16]. This functional, when regarded as a functional depending on the metric that determines the Dirac operator, is commonly known as the Einstein–Hilbert functional, since it directly corresponds to the Einstein–Hilbert action – the cornerstone of general relativity.
2 Spectral functionals: preliminaries
In this section, we first recall the results of computations of Wodzicki residues for a broad class of Laplace- and Dirac-type operators acting on sections of a vector bundle $V$ over a manifold $M$. Here, $M$ always denotes a compact, closed Riemannian manifold of dimension $n=2m>1$ with a fixed metric. [^1] We remark that most of the results presented here remain valid in the odd-dimensional case, as observed in [2] for the spin-Dirac operator. For the sake of concreteness and generalizations involving gradings, we focus here on even-dimensional manifolds.
We follow the conventions and notation used previously in [7, 2]; for example, we work in normal coordinates at a fixed point and expand the homogeneous symbols of the operators under consideration in these coordinates. We recall the notion of a Laplace-type operator discussed in [2]. An operator $L$ acting on $\Gamma(V)$, the module of smooth sections of $V$, is said to be of Laplace type if its symbol at any point $x\in M$ is given by $\sigma(L)=\mathfrak{a}_{2}+\mathfrak{a}_{1}+\mathfrak{a}_{0}$, where the homogeneous symbols $\mathfrak{a}_{\bullet}$, computed in normal coordinates $x$, are
$\mathfrak{a}_{2}=\left(\delta_{ab}+\frac{1}{3}R_{acbd}x^{c}x^{d}\right)\xi_{a}\xi_{b}+o\bigl({\bf x}^{2}\bigr),\qquad\mathfrak{a}_{1}={\rm i}\bigl(P_{ab}x^{b}+S_{a}\bigr)\xi_{a}+o({\bf x}),$
$\mathfrak{a}_{0}=Q+o({\bf 1}).$ (2.1)
Here, $R_{acbd}$ is the Riemann curvature tensor, and $P_{ab}$, $S_{a}$, and $Q$ are $\operatorname{End}(V)$-valued tensors evaluated at $x=0$. We recall that the Wodzicki residue [19] is (up to normalization) the unique trace on the algebra $\Psi\mathrm{DO}(V)$ of classical pseudodifferential operators
$\mathscr{W}\mathrm{res}(P)=\int_{M}\mathrm{dvol}_{g}\int_{|\xi|=1}{\rm d}^{4}x\,\mathrm{Tr}(\sigma_{-n}(P)(x,\xi)),$
where $\sigma_{-n}(P)$ denotes the symbol of $P\in\Psi\mathrm{DO}(V)$ of order n-n [10].
To compute the relevant functionals, we need the following general result.
Proposition 2.1 ([2, Proposition I.2]).
For a second-order differential operator $O$, with a symbol expressed in normal coordinates around a point on $M$,
$\sigma(O)=F^{ab}\xi_{a}\xi_{b}+{\rm i}G^{a}\xi_{a}+H+o({\bf 1}),$
where $F^{ab}=F^{ba}$, $G^{a}$, and $H$ are endomorphisms of the fiber $V$ at the point $x=0$, we have
$\mathscr{W}\mathrm{res}(OL^{-m})={}$ $\frac{\nu_{n-1}}{24}\int_{M}\mathrm{dvol}_{g}\mathrm{Tr}\bigl[24H+12G^{a}S_{a}+F^{aa}(-12Q+6P_{bb}-2R-3S_{b}S_{b})$
[…]
where $R$ is the scalar curvature and $\mathrm{Ric}$ is the Ricci tensor.
We now present two corollaries that will be particularly useful.
Corollary 2.2 ([2, Corollary I.3]).
For a $C^{\infty}(M)$-endomorphism $E\colon S\to S$, we have
$\mathscr{W}\mathrm{res}\bigl(EL^{-m+1}\bigr)=\frac{n-2}{24}\nu_{n-1}\int_{M}\mathrm{dvol}_{g}\mathrm{Tr}[E(-12Q+6P_{aa}-2R-3S_{a}S_{a})].$
Corollary 2.3 ([2, Corollary I.4]).
For an operator $O$ with a symbol as in Proposition 2.1, we have
$\mathscr{W}\mathrm{res}\left(\left(O-\frac{1}{n-2}F^{aa}L\right)L^{-m}\right)$
$\qquad{}=\frac{\nu_{n-1}}{24}\int_{M}\mathrm{dvol}_{g}\mathrm{Tr}[24H+12G^{a}S_{a}+2F^{ab}(-6P_{ab}+2\mathrm{Ric}_{ab}-3S_{a}S_{b})].$
We impose minimal assumptions on $D_{0}$ ensuring that $D_{0}^{2}$ is of Laplace type.
Definition 2.4.
We say that $D=D_{0}+B$ is a Dirac-type operator on sections of the vector bundle $V$ if the symbol of $D_{0}$ in normal coordinates around a fixed point on the manifold is given by $\sigma(D_{0})=\mathfrak{d}_{1}+\mathfrak{d}_{0}$, where
$\mathfrak{d}_{1}=\bigl(-\gamma^{a}+f^{a}_{bc}x^{b}x^{c}\bigr)\xi_{a}+o\bigl({\bf x}^{2}\bigr),\qquad\mathfrak{d}_{0}=g_{a}x^{a}+o({\bf x}),$
for some smooth endomorphisms $f^{a}_{bc}$ (symmetric in $b$, $c$) and $g_{a}$ of $V$ evaluated at $x=0$. Additionally, the endomorphism $B$ has an expansion $B=B_{0}+B_{a}x^{a}+o({\bf x})$.
The structure of $\mathfrak{d}_{1}$ and $\mathfrak{d}_{0}$, together with the requirement that $\mathfrak{d}_{1}^{2}=\mathfrak{a}_{2}$ as in (2.1), ensures that both $D_{0}^{2}$ and $D^{2}$ are Laplace-type. Notably, the absence of linear terms in $\bf x$} in the principal symbol $\mathfrak{d}_{1}$, and the vanishing of the zero-order symbol $\mathfrak{d}_{0}$ at $x=0$, are essential conditions [^2] One could define Dirac-type operators without requiring the absence of linear terms in $\bf x$} in the principal symbol. However, for our purposes, this condition is essential.. The explicit form of the symbols of $D_{0}$ (e.g., coefficients $f^{a}_{bc}$ and $g_{a}$) will not matter, as we focus only on the perturbation by $B$. The $\gamma^{a}$ are Clifford algebra generators on each fiber $V$ satisfying $\bigl\{\gamma^{a},\gamma^{b}\bigr\}=2\delta^{ab}\mathbf{1}$}. Next, only the first two terms in the expansion of $B$ are relevant, since $B$ does not enter the principal symbol of $D$. Finally, since $D^{2}=D_{0}^{2}+\{D_{0},B\}+B^{2}$, and using the relation $\{{\rm i}\gamma^{a}\partial_{a},B\}={\rm i}\{\gamma^{a},B\}\partial_{a}+{\rm i}\gamma^{a}\partial_{a}B$, we get
$P_{ab}=P_{ab}^{0}+{\rm i}\{\gamma^{a},B_{b}\},\qquad S_{a}={\rm i}\{\gamma^{a},B_{0}\},$
$S_{a}^{0}=0,\qquad Q=Q^{0}+{\rm i}\gamma^{a}B_{a}+B_{0}^{2},$ (2.2)
with $P^{0}_{ab}={\rm i}\{\gamma^{a},g_{b}\}+2\gamma^{c}f^{a}_{bc}$, $S_{a}^{0}=0$, $Q^{0}={\rm i}\gamma^{a}g_{a}$.
Proposition 2.5.
Let $D$ be as in Definition 2.4. Then, for $O=ED$, where $E$ is an arbitrary bundle endomorphism, we have
$\mathscr{W}\mathrm{res}\bigl(EDD^{-2m}\bigr)=\frac{\nu_{n-1}}{2}\int_{M}\mathrm{dvol}_{g}\mathrm{Tr}\left[E(2B_{0}-\gamma^{a}\{\gamma^{a},B_{0}\})\right].$
Proof.
Observe that $O=ED$ has the structure from Proposition 2.1, with $F^{ab}=0$, $G^{a}={\rm i}E\gamma^{a}$, and $H=EB_{0}$. The result follows directly from Proposition 2.1, with $L=D^{2}$, and the following observations: For $k>0$, the leading symbols of $D^{-2k}$ read $\sigma\bigl(D^{-2k}\bigr)=\mathfrak{c}_{2k}+\mathfrak{c}_{2k+1}+\mathfrak{c}_{2k+2}$, where
@@ 2 Spectral functionals: preliminaries
Here, $\mathfrak{c}_{p}^{0}$ denotes the respective symbols of $D_{0}^{-2k}$ $\sigma\bigl(D_{0}^{-2k}\bigr)=\mathfrak{c}_{2k}^{0}+\mathfrak{c}_{2k+1}^{0}+\mathfrak{c}_{2k+2}^{0}$. This, together with (2.2), concludes the claim. ∎
From now on, we assume that the Dirac operator fulfills the above requirements.